OpenStudy (anonymous):

How do i solve for x? 3^2x-1=3^x+2

OpenStudy (turingtest):

is this\[3^{2x}-1=3^x+2\]or\[3^{2x-1}=3^{x+2}\]?

Parth (parthkohli):

I guess that it is the 2nd one. When the base is same in the equations, we can for an equation with the exponents. \(\Large \color{midnightblue}{\rightarrow 2x - 1 = x + 2 }\) Why can we form an equation when the bases are same? Let's see, \(3^2 = 3^{x - 1}\) When we solve it, we get x as 3. Let's check. \(\Large \color{midnightblue}{\rightarrow 3^2 = 3^{3 - 1} }\) \(\Large \color{midnightblue}{\rightarrow 9 = {27 \over 3} }\) \(\Large \color{midnightblue}{\rightarrow 9 = 9 }\) Checked! \(\checkmark\) @TuringTest Check it please.

OpenStudy (turingtest):

that is indeed correct @ParthKohli provided that is the question

OpenStudy (anonymous):

turning test, its the second one

Parth (parthkohli):

Then my answer, see it.

OpenStudy (anonymous):

oh, k. Thanks

OpenStudy (anonymous):

wait i still dont get it